OpenStudy (anonymous):

Which of the following statements is NOT true? 1. If x+2/x=5+y/5, then 2/x=y/5 2. If x+3/4=y/5, then x+3/y=4/5 3. If x/3=y/4, then x/y=4/3 4. If x-4/4=3-y/y, then x/4=3/y

OpenStudy (anonymous):

if x/3 = y/4, then x/y = 4/3 is not true

OpenStudy (anonymous):

ok tyvm